# ReferencesSuppReferences

### v2

@@ Abstract
We provide precise conditions for nonparametric identification of causal effects by high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other shocks (separability) and (ii) the surprise component of the news or event dominates all other shocks that are present in the event window (relative exogeneity). Technically, the latter condition requires the policy shock to have infinite variance in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
Separability holds, for example, when the model is linear, which is often assumed in applied work. The key condition deserving of careful scrutiny in applications, relative exogeneity, holds when the policy shock has infinite variance while the other shocks have finite variance within the event window. The latter shocks correspond to factors that are not specific to the announcement and may also be present in non-announcement periods. In contrast, the policy shock occurs in a lumpy manner as the unexpected part of the news quickly spreads among economic agents. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can justify relative exogeneity. Even when the policy shock does not have infinite variance, which can be difficult to verify in practice, we show that the event study estimator has low bias for a weighted average of causal effects when the variance of the policy shock is large enough relative to that of other shocks in the window. In this sense, relative exogeneity can serve as a good approximation.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
and refer to this limiting case as relative exogeneity. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample so that for the other announcements, the policy news was not a surprise for the markets. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.
@@ 4.1 Response of Interest Rates to Monetary Policy News
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.232323We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the treatment sample. [^24] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.252525The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the treatment sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the treatment and control sample brackets the same period.
@@ 4.2 Response of Blue Chip Forecasts on Output to Monetary Policy News
This argument should also apply to the event study regression with Treasury yields as dependent variable. However, Bauer and Swanson (2023a) found that in regressions where the dependent variable is the change in an asset price or Treasury yield, controlling for macroeconomic and financial data that predate the FOMC announcement does not change the point estimates and their statistical significance. This difference likely arises because the changes in Treasury yields are constructed using a 30-minute or 1-day window around the announcement so that relative exogeneity is likely to provide a good approximation. As can be seen from Table 4, the changes in Treasury yields based on a 1-day window have an order of magnitude similar to those of the changes in the policy variable. Since relative exogeneity provides a good approximation in this case, controlling for $X_{t}$ does not result in a change in the point estimates even though $D_{t}$ and $X_{t}$ are correlated. The same omitted economic news that predate the announcement do not generate bias when $Y_{t}$ is the change of an asset price or Treasury yield over a similarly-sized narrow window used to construct the policy surprise. The key point here is that the validity of the event study approach does not require that there is no endogeneity. Rather, it requires that the policy shock dominates any other shock that is present in the event window.
@@
This discussion suggests that the event study approach is more credible when a narrow window is used for both the dependent and independent variable. Both 30-minute and 1-day windows are good choices as relative exogeneity is more likely to hold. If longer windows (e.g., one-month window) are used, as is the case for Blue Chip forecasts, then relative exogeneity is less likely to hold and the researcher should make more effort to appropriately control for omitted variables and reverse causality. For example, orthogonalizing surprises rather than using the original surprises as suggested by Bauer and Swanson (2023a) could provide a solution in these contexts.

### v3

@@ Abstract
We provide precise conditions for nonparametric identification of causal effects by high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other shocks (separability) and (ii) the surprise component of the news or event dominates all other shocks that are present in the event window (relative exogeneity). Technically, the latter condition requires the ratio between the variance of the policy shock and that of the other shocks to be infinite in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
Separability holds, for example, when the model is linear, which is often assumed in applied work. The key condition deserving of careful scrutiny in practice, relative exogeneity, holds when the policy shock has infinite variance while the other shocks have finite variance within the event window. More precisely, relative exogeneity holds when the ratio between the variance of the policy shock and that of the other shocks (i.e., background noise) is infinite in the event window. Thus, relative exogeneity also holds when the variance of the policy shock is finite while the variance of the background noise is vanishing. The latter shocks correspond to factors that are not specific to the announcement and may also be present in non-announcement periods. In contrast, the policy shock occurs in a lumpy manner as the unexpected part of the news quickly spreads among economic agents. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can justify relative exogeneity. Even when the policy shock does not have infinite variance, which can be difficult to verify in practice, we show that the event study estimator has low bias for a weighted average of causal effects when the variance of the policy shock is large enough relative to that of other shocks in the window. In this sense, relative exogeneity can be seen as an idealized limiting case that can serve as a good approximation to the practically-relevant case of a very large variance ratio for the policy shock relative to the other shocks in the window.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
and refer to this limiting case as relative exogeneity. [^10] More precisely, relative exogeneity holds when the ratios between the variance of the policy shock and the variances of the other shocks in the event window diverge. Thus, one may instead frame relative exogeneity as the condition for which the variance of the policy shock is finite and the variance of the background noise vanishes in the event window. The same identification results can also be shown to hold under this latter condition. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample so that for the other announcements, the policy news was not a surprise for the markets. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.
@@ 2.3 Identification Results
Note that the high-frequency event study method is different from heteroskedasticity-based identification [cf. rigobon:2003], sometimes referred to as Rigobon’s method. The latter also uses information from the control sample, which includes windows (either 30-minutes or 1-day windows) that do not bracket an announcement or event. The main heteroskedasticity-based identification condition is that the volatility of the policy shock is larger in the event windows than in the control windows. This is different from relative exogeneity for two reasons. First, the required increase in volatility of the policy shock is not relative to the other shocks in the event window but relative to the policy shock in the control windows. Second, the increase in volatility need not be infinite. Thus, heteroskedasticity-based identification requires neither stronger nor weaker conditions than the relative exogeneity condition. In addition, Rigobon’s estimator is an instrumental variables estimator and therefore different from the OLS high-frequency event study estimator. Since these two estimators correspond to different estimands, they identify different causal effect in general.
@@ 4.1 Response of Interest Rates to Monetary Policy News
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.242424We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the treatment sample. [^25] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.262626The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the treatment sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the treatment and control sample brackets the same period.
@@ 4.2 Response of Blue Chip Forecasts on Output to Monetary Policy News
This argument should also apply to the event study regression with Treasury yields as dependent variable. However, Bauer and Swanson (2023a) found that in regressions where the dependent variable is the change in an asset price or Treasury yield, controlling for macroeconomic and financial data that predate the FOMC announcement does not change the point estimates and their statistical significance. This difference likely arises because the changes in Treasury yields are constructed using a 30-minute or 1-day window around the announcement so that relative exogeneity is likely to provide a good approximation. As can be seen from Table 4, the changes in Treasury yields based on a 1-day window have an order of magnitude similar to those of the changes in the policy variable. Since relative exogeneity provides a good approximation in this case, controlling for $X_{t}$ does not result in a change in the point estimates even though $D_{t}$ and $X_{t}$ are correlated. The same omitted economic news that predate the announcement do not generate bias when $Y_{t}$ is the change of an asset price or Treasury yield over a similarly-sized narrow window used to construct the policy surprise. The key point here is that the validity of the event study approach does not require the absence of endogeneity. Rather, it requires that the policy shock dominates any other shock that is present in the event window. When this condition holds, identification of causal effects from an event study does not require the inclusion of controls in the event study regression. It is interesting to note that this implies that, when the identification conditions are met, event study regressions are immune to forms of p-hacking that involve searching through different control specifications, a potential strength of the high-frequency event study method.
This discussion suggests that the event study approach is more credible when a narrow window is used for both the dependent and independent variable. Both 30-minute and 1-day windows are good choices as relative exogeneity is more likely to hold. This was also informally discussed on p. 1289 in Nakamura and Steinsson (2018a). If longer windows (e.g., one-month window) are used, as is the case for Blue Chip forecasts, then relative exogeneity is less likely to hold and the researcher should make more effort to appropriately control for omitted variables and reverse causality. For example, orthogonalizing surprises rather than using the original surprises as suggested by Bauer and Swanson (2023a) could provide a solution in these contexts.